# Singularity-free cosmology from interactions in the dark sector

### v1

@@ V Dynamics with $q>3$
Although non-singular models exist in the $q<3$ phase spaces, our main result lies in the $q>3$ case, as all trajectories are non-singular for any initial condition. As before, we show the $Z$-$X$ phase spaces, and take $Y>0$ ($H>0$) so the trajectories are expanding. Again, we set $\mathcal{R}=0.01$ to produce readable phase spaces. As in the $q<3$ cases, there are three possible types of stability character for the $dS_{1+}$ fixed point: a repelling node, an improper node or a spiral repellor. In the 2-D phase spaces, only one separatrix exists, which is the repellor-saddle separatrix that joins the $dS_{1+}$ repellor and the $dS_{3+}$ saddle (see Fig. 1). When $q>3$, all the critical points representing singularities have generalized saddle stability, which means all trajectories avoid the singularities, and expand from $dS_{1+}$ and asymptotically tend towards $dS_{2+}$ at late times. The zero acceleration curve either crosses the $Z=0$ axes below the $dS_{2+}$ attractor, or does not cross it at all, therefore all trajectories have late-time acceleration as they approach a cosmological constant. The behavior of the trajectories at low energy is as in Fig. 2(b).
As before, the repellor-saddle separatrix may or may not intersect the zero acceleration ($a^{\prime\prime}=0$) curve. When $dS_{1+}$ is a spiral repellor, three subcases exist for the 2-D phase space: the repellor-saddle separatrix and $a^{\prime\prime}=0$ curve can intersect twice, touch or do not intersect at all depending on the specific value of $q$. However, when $dS_{1+}$ is a repelling or an improper node, the only subcase that exists is when the separatrix and zero acceleration curve do not intersect. We do not focus on the cases where $dS_{1+}$ is an improper or repelling node here, however we explore them in the Appendix and highlight the physically interesting models.
The subcase where $dS_{1+}$ is a spiral repellor, and the repellor-saddle separatrix intersects the $a^{\prime\prime}=0$ curve twice is particularly important, as all trajectories in this phase space are qualitatively interesting. In this case, trajectories can intersect the zero acceleration curve two, three or four times, which correspond to first integral surfaces in 3-D which have two, three or four Einstein fixed points, respectively. In the following, we present the $Z$-$X$ phase spaces where $dS_{1+}$ is a spiral repellor, so we can show all possible models. The color scheme of the fixed points and critical points, and the color scheme of the curves are given in Tables 4 and 5, respectively.
@@ V.1.1 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve do not intersect
Figure 7: The $Z$-$X$ phase space referred to in Sec. V.1.1. Here, the parameters are set to $w_{x}=-0.8$ and $q=3.5$. All trajectories expand from $dS_{1+}$ (orange) and evolve to $dS_{2+}$ (magenta), except for the separatrix which expands to $dS_{3+}$ (dark blue). Therefore, all models avoid the singularities. In this case, however, the separatrix does not intersect the zero acceleration curve. Consequently, there are trajectories that always accelerate, and there are trajectories which intersect the $a^{\prime\prime}=0$ curve twice and have a decelerated period. There is also a trajectory that touches the zero acceleration curve, but otherwise accelerates.
The first subcase is when the repellor-saddle separatrix and $a^{\prime\prime}=0$ curve do not intersect, which is shown in Fig. 7. Not all trajectories intersect the $a^{\prime\prime}=0$ curve, and those trajectories correspond to first integral surfaces in 3-D that have no Einstein fixed points. These trajectories never have a decelerated period, and so are not of interest. A trajectory exists that touches, but does not intersect the $a^{\prime\prime}=0$ curve. This trajectory corresponds to a surface in 3-D that has one Einstein fixed point, and has a point where there is no acceleration, but never decelerates so is not of interest. Finally, there are trajectories that intersect the $a^{\prime\prime}=0$ curve twice. These correspond to surfaces in the 3-D phase space with two Einstein fixed points. Initially these trajectories accelerate, then decelerate when they cross below the red $a^{\prime\prime}=0$ curve, and then cross back above the $a^{\prime\prime}=0$ curve and accelerate as they asymptotically approach $dS_{2+}$.
@@ V.1.2 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve touch
Fig. 8 shows the subcase where the separatrix and $a^{\prime\prime}=0$ curve touch. In this case, the separatrix itself is a trajectory which corresponds to a surface in the $X$-$Y$-$Z$ phase space with one Einstein fixed point. This trajectory reaches a point with zero acceleration, however it never decelerates. All other trajectories intersect the $a^{\prime\prime}=0$ curve twice, initially accelerating, then decelerating when they cross below the $a^{\prime\prime}=0$ curve, and accelerating again at late times as they asymptotically expand toward $dS_{2+}$. These trajectories have corresponding first integral surfaces in 3-D that have two Einstein fixed points.
Figure 8: The $Z$-$X$ phase space referred to in V.1.2. Here, the parameters are set to $w_{x}=-0.9$ and $q=4.2$. In this case, the separatrix touches the zero acceleration curve, therefore the separatrix itself is a trajectory that reaches a point of zero acceleration, but otherwise accelerates. All other trajectories initially accelerate as they expand from $dS_{1+}$, then cross below the red $a^{\prime\prime}=0$ curve and have a decelerated period, and finally cross above the $a^{\prime\prime}=0$ curve again, accelerating as they evolve to $dS_{2+}$.
@@ V.1.3 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve intersect twice
Figure 9: The $Z$-$X$ phase space referred to in V.1.3. Here, the parameters are set to $w_{x}=-0.95$ and $q=3.1$. Fig 9(a) shows the full phase space, where all trajectories expand from $dS_{1+}$ (orange) and evolve to $dS_{2+}$ (magenta), except for the saddle-repellor separatrix which expands to $dS_{3+}$ (dark blue). The separatrix intersects the zero acceleration curve twice, which means all trajectories have a decelerated period. Trajectories which intersect the zero acceleration curve four times have two periods of deceleration, and those that intersect it twice have one decelerated phase. There is a trajectory that expands from the repellor and touches the zero acceleration curve and then intersects it twice. All trajectories in this case are of interest, as all are non-singular and have a decelerated phase where large-scale structure could form. Fig 9(b) shows the behavior of the trajectories as they approach $dS_{2+}$. Some trajectories have decreasing dark energy density and evolve towards $dS_{2+}$ from above the green $X^{\prime}=0$ curve, and some cross below the $X^{\prime}=0$ curve and exhibit phantom behavior with increasing dark energy density.
@@
The final subcase is where the separatrix and zero acceleration curve intersect twice, which is shown in Fig. 9. In this case all trajectories have at least one period of deceleration, and are therefore all of interest. Some trajectories intersect the $a^{\prime\prime}=0$ curve twice, which correspond to surfaces in the $X$-$Y$-$Z$ phase space that have two Einstein fixed points. These trajectories initially accelerate, then have one period of deceleration, and accelerate again at late times. One trajectory in the phase space will touch the $a^{\prime\prime}=0$ curve, and then intersect it twice; this trajectory also has one decelerated period, and corresponds to a surface in 3-D that has three Einstein fixed points. Finally, there are trajectories which intersect the $a^{\prime\prime}=0$ curve four times, which correspond to first integral surfaces in 3-D that have four Einstein fixed points. These trajectories have two periods of deceleration when they cross below the $a^{\prime\prime}=0$ curve, and have three acceleration periods when they are above it. All of the trajectories here are of interest, therefore we concentrate our analysis in 3-D on this subcase.
V.2 3-D Phase Spaces
@@
We will focus our analysis in 3-D on the subcase shown in Fig. 9, as all of these trajectories qualitatively match observations, as they have a decelerated matter-dominated era followed by a late-time acceleration. There are three cases we present in the following sections, as the trajectories in Fig. 9 can correspond to first integral surfaces in 3-D that have two, three or four Einstein fixed points. Some of the trajectories plotted in Fig.s 12 - 16 have orange sections. The numerical solutions in Mathematica break down close to $S_{2\pm}$, so it is not possible to plot some trajectories completely numerically. In order to plot these trajectories fully, we patch together the numeric solutions with analytic approximations. We solve the equations numerically until the solution breaks down. We then take initial conditions from the numeric solution, and solve the high energy analytic approximation in Eq. (29) around the singularity, which we plot in orange. We then take initial conditions from the high energy analytic approximation to plot the rest of the trajectory numerically. The color scheme of the fixed points and critical points in the 3-D phase spaces are given in Table 8, and the color scheme of the curves and surfaces are given in Table 9.
@@ V.2.1 Two Einstein Fixed Points
Fig.s 10 and 11 show a first integral surface in the $X$-$Y$-$Z$ phase space with two Einstein fixed points, which corresponds to a trajectory in Fig. 9 that intersects the $a^{\prime\prime}=0$ curve twice. One Einstein point is a center, and the other is a saddle which the CFS passes through. Fig. 10 shows only the trajectories with positive spatial curvature, which all have a bounce. Trajectories that evolve outside of the CFS, and a subset that evolve inside the CFS, bounce once and contract from $dS_{2-}$ and expand to $dS_{2+}$. There is also a subset of trajectories that are cyclic around the center Einstein fixed point. These repeatedly expand, turnaround, contract and bounce. The trajectories inside the CFS are not of interest, as the bouncing models always accelerate, and the cyclic models have an initial accelerated expansion followed by a decelerated period, but they turn around and contract during the decelerated period so there is no late-time accelerated expansion. As for the $q<3$ case, the trajectories that bounce once outside of the CFS are the models of interest. They have an accelerated era in which they evolve toward a high energy quasi-de Sitter phase when they bounce, and then become flatter as they start to expand; they then have a decelerated expanding phase followed by a final period of acceleration as they evolve toward the low energy cosmological constant represented by $dS_{2+}$.
@@
Figure 10: The $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.02$ and $Z_{0}=0.2$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$ and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$. During expansion, these trajectories cross the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
Fig. 11 shows the expanding flat and open trajectories which are qualitatively interesting. These trajectories emerge from the de Sitter fixed points $dS_{1+}$ and $dS_{4+}$, respectively, and cross the zero acceleration curve twice, so have a decelerated period followed by a final late-time acceleration as they asymptotically approach $dS_{2+}$.
@@
Figure 11: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.02$ and $Z_{0}=0.2$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve twice. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and cross the zero acceleration surface twice, meaning they initially accelerate, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ V.2.2 Three Einstein Fixed Points
Fig.s 12 - 14 show a first integral surface in 3-D with three Einstein fixed points, which corresponds to the trajectory in Fig. 9 which touches the $a^{\prime\prime}=0$ curve, and then intersects it twice. Fig. 12 shows the trajectories with positive spatial curvature. A CFS exists through the cusp Einstein fixed point in the middle, which is plotted in the figure. The trajectories above and below this CFS all contract from $dS_{2-}$, then accelerate through a bounce and become flatter as they start to expand; the trajectories above the CFS bounce during a quasi-de Sitter phase. The bouncing models then have a decelerated expanding period followed by a late-time acceleration as they evolve towards $dS_{2+}$. We expect another CFS to exist through the Einstein fixed point close to $X=0$, which is a saddle, and loop around the Einstein fixed point close to $X=1$, which is a center (similar to the CFS in Fig. 10). We have not plotted this separatrix due to the numerics breaking down at high energy, however we do not expect trajectories within this CFS to be qualitatively interesting; we expect there are bouncing trajectories that always accelerate, and cyclic trajectories which have no late-time acceleration. Fig. 13 shows the side view of the 3-D phase space through the $Z$-$X$ plane, which more clearly shows the corresponding trajectory in Fig. 9.
@@
Figure 12: The $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.145$ and $Z_{0}=0.1$, which corresponds to a trajectory in Fig. 9 that first touches the red zero acceleration curve, and then intersects it twice. In this figure, only the trajectories with positive spatial curvature are plotted. Three Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. All trajectories bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$, and all cross the red zero acceleration surface twice, so have a decelerated phase followed by a late-time acceleration.
@@
Figure 13: The side view of Fig. 12. All trajectories intersect the red zero acceleration surface twice during expansion, accelerating when they are above the surface and decelerating when they are below it.
@@
Fig. 14 shows the expanding flat and open trajectories, which emerge from the de Sitter fixed points $dS_{1+}$ and $dS_{4+}$, respectively, and asymptotically approach $dS_{2+}$ at late-times. Both flat and open models intersect the $a^{\prime\prime}=0$ curve twice, so have a decelerated period followed by a final late-time acceleration.
@@
Figure 14: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.145$ and $Z_{0}=0.1$, which corresponds to a trajectory in Fig. 9 that first touches the red zero acceleration curve, and then intersects it twice. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and then touch the red zero acceleration surface before crossing it twice, eventually expanding toward $dS_{2+}$ (magenta) at low energy.

### v2

@@ V Dynamics with $q>3$
Although nonsingular models exist in the $q<3$ phase spaces, our main result lies in the $q>3$ case, as all trajectories are nonsingular for any initial condition. As before, we show the $Z$-$X$ phase spaces, and take $Y>0$ ($H>0$) so the trajectories are expanding. Again, we set $\mathcal{R}=0.01$ to produce readable phase spaces. As in the $q<3$ cases, there are three possible types of stability character for the $dS_{1+}$ fixed point: a repelling node, an improper node or a spiral repellor. In the two-dimensional phase spaces, only one separatrix exists, which is the repellor-saddle separatrix that joins the $dS_{1+}$ repellor and the $dS_{3+}$ saddle (see Fig. 1). When $q>3$, all the critical points representing singularities have generalized saddle stability, which means all trajectories avoid the singularities, and expand from $dS_{1+}$ and asymptotically tend towards $dS_{2+}$ at late times. The zero acceleration curve either crosses the $Z=0$ axes below the $dS_{2+}$ attractor, or does not cross it at all, therefore all trajectories have late-time acceleration as they approach a cosmological constant. The behavior of the trajectories at low energy is as in Fig. 2(2(b)).
As before, the repellor-saddle separatrix may or may not intersect the zero acceleration ($a^{\prime\prime}=0$) curve. When $dS_{1+}$ is a spiral repellor, three subcases exist for the two-dimensional phase space: the repellor-saddle separatrix and $a^{\prime\prime}=0$ curve can intersect twice, touch or do not intersect at all depending on the specific value of $q$. However, when $dS_{1+}$ is a repelling or an improper node, the only subcase that exists is when the separatrix and zero acceleration curve do not intersect. We do not focus on the cases where $dS_{1+}$ is an improper or repelling node here; however, we explore them in the Appendix and highlight the physically interesting models.
The subcase where $dS_{1+}$ is a spiral repellor, and the repellor-saddle separatrix intersects the $a^{\prime\prime}=0$ curve twice is particularly important, as all trajectories in this phase space are qualitatively interesting. In this case, trajectories can intersect the zero acceleration curve two, three or four times, which correspond to first integral surfaces in three dimensions which have two, three or four Einstein fixed points, respectively. In the following, we present the $Z$-$X$ phase spaces where $dS_{1+}$ is a spiral repellor, so we can show all possible models. The color scheme of the fixed points and critical points, and the color scheme of the curves are given in Tables 4 and 5, respectively.
@@ V.1.1 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve do not intersect
Figure 7: $Z$-$X$ phase space referred to in Sec. V.1.1. Here, the parameters are set to $w_{x}=-0.8$ and $q=3.5$. All trajectories expand from $dS_{1+}$ (orange) and evolve to $dS_{2+}$ (magenta), except for the separatrix which expands to $dS_{3+}$ (dark blue). Therefore, all models avoid the singularities. In this case; however,, the separatrix does not intersect the zero acceleration curve. Consequently, there are trajectories that always accelerate, and there are trajectories which intersect the $a^{\prime\prime}=0$ curve twice and have a decelerated period. There is also a trajectory that touches the zero acceleration curve, but otherwise accelerates.
The first subcase is when the repellor-saddle separatrix and $a^{\prime\prime}=0$ curve do not intersect, which is shown in Fig. 7. Not all trajectories intersect the $a^{\prime\prime}=0$ curve, and those trajectories correspond to first integral surfaces in three dimensions that have no Einstein fixed points. These trajectories never have a decelerated period, and so are not of interest. A trajectory exists that touches, but does not intersect the $a^{\prime\prime}=0$ curve. This trajectory corresponds to a surface in three dimensions that has one Einstein fixed point, and has a point where there is no acceleration, but never decelerates so is not of interest. Finally, there are trajectories that intersect the $a^{\prime\prime}=0$ curve twice. These correspond to surfaces in the three-dimensional phase space with two Einstein fixed points. Initially these trajectories accelerate, then decelerate when they cross below the red $a^{\prime\prime}=0$ curve, and then cross back above the $a^{\prime\prime}=0$ curve and accelerate as they asymptotically approach $dS_{2+}$.
@@ V.1.2 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve touch
Figure 8 shows the subcase where the separatrix and $a^{\prime\prime}=0$ curve touch. In this case, the separatrix itself is a trajectory which corresponds to a surface in the $X$-$Y$-$Z$ phase space with one Einstein fixed point. This trajectory reaches a point with zero acceleration; however, it never decelerates. All other trajectories intersect the $a^{\prime\prime}=0$ curve twice, initially accelerating, then decelerating when they cross below the $a^{\prime\prime}=0$ curve, and accelerating again at late times as they asymptotically expand toward $dS_{2+}$. These trajectories have corresponding first integral surfaces in three dimensions that have two Einstein fixed points.
Figure 8: $Z$-$X$ phase space referred to in Sec. V.1.2. Here, the parameters are set to $w_{x}=-0.9$ and $q=4.2$. In this case, the separatrix touches the zero acceleration curve, therefore the separatrix itself is a trajectory that reaches a point of zero acceleration, but otherwise accelerates. All other trajectories initially accelerate as they expand from $dS_{1+}$, then cross below the red $a^{\prime\prime}=0$ curve and have a decelerated period, and finally cross above the $a^{\prime\prime}=0$ curve again, accelerating as they evolve to $dS_{2+}$.
@@ V.1.3 The repellor-saddle separatrix and $a^{\prime\prime}=0$ curve intersect twice
Figure 9: $Z$-$X$ phase space referred to in Sec. V.1.3. Here, the parameters are set to $w_{x}=-0.95$ and $q=3.1$. (9(a)) shows the full phase space, where all trajectories expand from $dS_{1+}$ (orange) and evolve to $dS_{2+}$ (magenta), except for the saddle-repellor separatrix which expands to $dS_{3+}$ (dark blue). The separatrix intersects the zero acceleration curve twice, which means all trajectories have a decelerated period. Trajectories which intersect the zero acceleration curve four times have two periods of deceleration, and those that intersect it twice have one decelerated phase. There is a trajectory that expands from the repellor and touches the zero acceleration curve and then intersects it twice. All trajectories in this case are of interest, as all are nonsingular and have a decelerated phase where large-scale structure could form. (9(b)) shows the behavior of the trajectories as they approach $dS_{2+}$. Some trajectories have decreasing dark energy density and evolve towards $dS_{2+}$ from above the green $X^{\prime}=0$ curve, and some cross below the $X^{\prime}=0$ curve and exhibit phantom behavior with increasing dark energy density.
@@
The final subcase is where the separatrix and zero acceleration curve intersect twice, which is shown in Fig. 9. In this case all trajectories have at least one period of deceleration, and are therefore all of interest. Some trajectories intersect the $a^{\prime\prime}=0$ curve twice, which correspond to surfaces in the $X$-$Y$-$Z$ phase space that have two Einstein fixed points. These trajectories initially accelerate, then have one period of deceleration, and accelerate again at late times. One trajectory in the phase space will touch the $a^{\prime\prime}=0$ curve, and then intersect it twice; this trajectory also has one decelerated period, and corresponds to a surface in three dimensions that has three Einstein fixed points. Finally, there are trajectories which intersect the $a^{\prime\prime}=0$ curve four times, which correspond to first integral surfaces in three dimensions that have four Einstein fixed points. These trajectories have two periods of deceleration when they cross below the $a^{\prime\prime}=0$ curve, and have three acceleration periods when they are above it. All of the trajectories here are of interest, therefore we concentrate our analysis in three dimensions on this subcase.
V.2 Three-dimensional phase spaces
@@
We will focus our analysis in three dimensions on the subcase shown in Fig. 9, as all of these trajectories qualitatively match observations, as they have a decelerated matter dominated era followed by a late-time acceleration. There are three cases we present in the following sections, as the trajectories in Fig. 9 can correspond to first integral surfaces in three dimensions that have two, three or four Einstein fixed points. Some of the trajectories plotted in Figs. 1216 have orange sections. The numerical solutions in Mathematica break down close to $S_{2\pm}$, so it is not possible to plot some trajectories completely numerically. In order to plot these trajectories fully, we patch together the numeric solutions with analytic approximations. We solve the equations numerically until the solution breaks down. We then take initial conditions from the numeric solution, and solve the high energy analytic approximation in Eq. (30) around the singularity, which we plot in orange. We then take initial conditions from the high energy analytic approximation to plot the rest of the trajectory numerically. The color scheme of the fixed points and critical points in the three-dimensional phase spaces are given in Table 8, and the color scheme of the curves and surfaces are given in Table 9.
@@ V.2.1 Two Einstein fixed points
Figures 10 and 11 show a first integral surface in the $X$-$Y$-$Z$ phase space with two Einstein fixed points, which corresponds to a trajectory in Fig. 9 that intersects the $a^{\prime\prime}=0$ curve twice. One Einstein point is a center, and the other is a saddle which the CFS passes through. Figure 10 shows only the trajectories with positive spatial curvature, which all have a bounce. Trajectories that evolve outside of the CFS, and a subset that evolve inside the CFS, bounce once and contract from $dS_{2-}$ and expand to $dS_{2+}$. There is also a subset of trajectories that are cyclic around the center Einstein fixed point. These repeatedly expand, turnaround, contract and bounce. The trajectories inside the CFS are not of interest, as the bouncing models always accelerate, and the cyclic models have an initial accelerated expansion followed by a decelerated period, but they turn around and contract during the decelerated period so there is no late-time accelerated expansion. As for the $q<3$ case, the trajectories that bounce once outside of the CFS are the models of interest. They have an accelerated era in which they evolve toward a high energy quasi-de Sitter phase when they bounce, and then become flatter as they start to expand; they then have a decelerated expanding phase followed by a final period of acceleration as they evolve toward the low energy cosmological constant represented by $dS_{2+}$.
@@
Figure 10: $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.02$ and $Z_{0}=0.2$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$ and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$. During expansion, these trajectories cross the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
Figure 11 shows the expanding flat and open trajectories which are qualitatively interesting. These trajectories emerge from the de Sitter fixed points $dS_{1+}$ and $dS_{4+}$, respectively, and cross the zero acceleration curve twice, so have a decelerated period followed by a final late-time acceleration as they asymptotically approach $dS_{2+}$.
@@
Figure 11: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.02$ and $Z_{0}=0.2$, which corresponds to a trajectory in Fig. 9 that intersects the red zero acceleration curve twice. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and cross the zero acceleration surface twice, meaning they initially accelerate, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ V.2.2 Three Einstein fixed points
Figures 1214 show a first integral surface in three dimensions with three Einstein fixed points, which corresponds to the trajectory in Fig. 9 which touches the $a^{\prime\prime}=0$ curve, and then intersects it twice. Figure 12 shows the trajectories with positive spatial curvature. A CFS exists through the cusp Einstein fixed point in the middle, which is plotted in the figure. The trajectories above and below this CFS all contract from $dS_{2-}$, then accelerate through a bounce and become flatter as they start to expand; the trajectories above the CFS bounce during a quasi-de Sitter phase. The bouncing models then have a decelerated expanding period followed by a late-time acceleration as they evolve towards $dS_{2+}$. We expect another CFS to exist through the Einstein fixed point close to $X=0$, which is a saddle, and loop around the Einstein fixed point close to $X=1$, which is a center (similar to the CFS in Fig. 10). We have not plotted this separatrix due to the numerics breaking down at high energy; however, we do not expect trajectories within this CFS to be qualitatively interesting; we expect there are bouncing trajectories that always accelerate, and cyclic trajectories which have no late-time acceleration. Figure 13 shows the side view of the three-dimensional phase space through the $Z$-$X$ plane, which more clearly shows the corresponding trajectory in Fig. 9.
@@
Figure 12: $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.145$ and $Z_{0}=0.1$, which corresponds to a trajectory in Fig. 9 that first touches the red zero acceleration curve, and then intersects it twice. In this figure, only the trajectories with positive spatial curvature are plotted. Three Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. All trajectories bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$, and all cross the red zero acceleration surface twice, so have a decelerated phase followed by a late-time acceleration.
@@
Figure 13: Side view of Fig. 12. All trajectories intersect the red zero acceleration surface twice during expansion, accelerating when they are above the surface and decelerating when they are below it.
@@
Figure 14 shows the expanding flat and open trajectories, which emerge from the de Sitter fixed points $dS_{1+}$ and $dS_{4+}$, respectively, and asymptotically approach $dS_{2+}$ at late times. Both flat and open models intersect the $a^{\prime\prime}=0$ curve twice, so have a decelerated period followed by a final late-time acceleration.
@@
Figure 14: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 9, where $w_{x}=-0.95$ and $q=3.1$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.145$ and $Z_{0}=0.1$, which corresponds to a trajectory in Fig. 9 that first touches the red zero acceleration curve, and then intersects it twice. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and then touch the red zero acceleration surface before crossing it twice, eventually expanding toward $dS_{2+}$ (magenta) at low energy.